OpenStudy (anonymous):

The box is to have a square base and hold 32in^ 3. Find the dimensions of the box that can be built with the minimum amount of material. You must justify your answer.

OpenStudy (aum):

Does it say whether the box is closed or open?

OpenStudy (aum):

Assume the square base is x inches by x inches and the height is h. Volume of the box = x * x * h = x^2h = 32 ---- (1) "minimum amount of material" implies minimum surface area. Assume the box is open, the surface area A of the box is: A = x^2 + 4xh ----- (2) From (1): h = 32 / x^2 From (2): A = x^2 + 4x * 32 / x^2 A = x^2 + 128/x = x^2 + 128x^(-1) Minimize A by finding the derivative and equating it to zero. dA/dx = 2x -128x^(-2) = 0 2x - 128/x^2 = 0 multiply by x^2: 2x^3 - 128 = 0 2x^3 = 128 x^3 = 64 x = 4 inches Put it in (1): 4^2 * h = 32 16h = 32 h = 2 The dimensions of the box are: 4 in. x 4 in. x 2 in. PS: If it is a closed box the we have to include the surface area of the lid and equation (2) will become A = 2x^2 + 4xh which you can minimize as before.